simplify the rational expresion
6 - 2x
-----
4x - 12

Answer by stanbon(48558)

(6 - 2x)/(4x - 12)
=2(3-x)/4(x-3)
Notice that 3-x = -(x-3)
= -1/2
Cheers,
Stan H.